OpenStudy (rockinhood):

I will give a medal for the best answer (obviously) For the data in the table, tell whether y varies directly with x. x y -6 9 1 -1.5 8 -12

OpenStudy (anonymous):

y varies directly with x (or: "x and y are directly proportional") if there is a constant k such that y = kx 9 = k(-6) -1.5 = k(1) 12 = k(8) If y varies directly with x, then k must equal -1.5 9 = (-1.5)(6) = 9 true. 12 = (-1.5)(8) = 12 true. So yes y varies directly with x.